OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. c√c + √7x - 2c√c

OpenStudy (anonymous):

we see there is a like term of csqrtc so we get -c(sqrtc) + sqrt7x

OpenStudy (saifoo.khan):

We always add like terms.

OpenStudy (saifoo.khan):

Same as in rational expressions with variables of different powers.

OpenStudy (anonymous):

It gives me only 4 choices though... 11c, c√c, 11c√c, or 11√c.

OpenStudy (anonymous):

I agree with Saif because I don't know what to do otherwise.

OpenStudy (saifoo.khan):

Is your question correct?

OpenStudy (anonymous):

Yup, sure is. It's probably the school.

OpenStudy (anonymous):

What about this one? Find the exact value of the radical expression in simplest form. √169 - √50 - √8

OpenStudy (saifoo.khan):

For this one, we have to factor the terms out.

OpenStudy (saifoo.khan):

i mean factor the square roots .

OpenStudy (saifoo.khan):

Lol, dont sleep in class! ;D

OpenStudy (anonymous):

I don't. I'm actually really dedicated. Algebra is just really hard for me.

OpenStudy (saifoo.khan):

http://www.youtube.com/watch?v=6QJtWfIiyZo

OpenStudy (saifoo.khan):

@fallingangel

OpenStudy (anonymous):

Thanks!

OpenStudy (saifoo.khan):

I can't do any better than that guy^

OpenStudy (anonymous):

guy? :P well, i appreciate it anyway.

OpenStudy (saifoo.khan):

The guy in the vid @fallingangel

OpenStudy (anonymous):

√169 - √50 - √8 for this radical we know √169 = 13, √50 = √2*25 =√2*√25= 5*√2 and for √8 = √4*2 = √4*√2 = 2√2 so now let us plug this in so we got 13 -5√2 - 2√2 = 13 -7√2 :)

OpenStudy (anonymous):

OMIGOSH! thank you! <3

OpenStudy (anonymous):

@hamza_b23

OpenStudy (anonymous):

lol no worries

OpenStudy (anonymous):

so just became became your fan... haha

OpenStudy (anonymous):

soo i see :P thanks :)